OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 8 - 40 + 200 - 1000 + ...

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

Each term differs by a factor of -5: \[8(5^0)=8\\ 8(5^1)=40\\ 8(5^2)=200\\\vdots\]Also, each successive term changes sign. So, using sigma notation, you have \[\sum_{n=0}^\infty8(-1)^n(5)^n=\sum_{n=0}^\infty8(-5)^n\]

OpenStudy (anonymous):

Thank you so very much!